OpenStudy (anonymous):

graph the equation by plotting three points 2y=5x+11

OpenStudy (anonymous):

(1,8) (3,13) (-1,3)

OpenStudy (anonymous):

3,13 is not one of the answers ibsmart

OpenStudy (anonymous):

that's not the answer

OpenStudy (zale101):

oh wait, you're trying to find the three points not the graph

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do you mean answer? If you substitute 13 for y and 3 for x into the equation, it works.

OpenStudy (anonymous):

13 is not one of the points on the graphg

OpenStudy (anonymous):

I need to find what the three points that make the equation true

OpenStudy (campbell_st):

use a table of values x: -1 : 1 : 3 substitute them into the equation to find y --------------- y: 3 : 8 : 13 plot the ordered pairs, (-1, 3) (1, 8) and (3, 13) on the numberplane then draw a long line through all 3 points

OpenStudy (anonymous):

2(13)=5(3)+11 26=15+11 26=26 True

OpenStudy (anonymous):

the point 13 is not on the graph it goes up to 10

OpenStudy (campbell_st):

why not just substitute an x value to find a y value and the ordered pair will be on the line x = 1 2y = 5*1 + 11 2y = 16 so y = 8 plot the ordered pair (1, 8) repeat the process with values like -1 and 3

OpenStudy (anonymous):

Then try (-3,-2)

OpenStudy (campbell_st):

well it is on the line, its just that the number plane is not large enough to include it so choose x = -3, x = -1 and x = 1 substitute those values to find the ordered pairs

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

in a question of this type.... you choose the values for x.... then substitute to find the corresponding y values if you do it correctly you can draw a line through all 3 points

OpenStudy (anonymous):

I got (-3,-4) (-1,6) (1,16)

OpenStudy (campbell_st):

how can you get an ordered pair of x = -1 and y = 6 \[2y = 5 \times (-1) + 11\] then \[2y = -5 + 11\] so \[2y = 6\] what is y...?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

can someone please help me

OpenStudy (campbell_st):

so the 1st point to plot on the number plane is (-1, 3) now for x = 1 \[2y = 5 \times 1 + 11\] I'll let you solve this

OpenStudy (anonymous):

1,16

OpenStudy (campbell_st):

no same mistake as before 2y = 16 what is y..?

OpenStudy (anonymous):

8

OpenStudy (campbell_st):

so plot (1, 8) now substitute x = -3 \[2y = 5 \times (-3) + 11\] whats the value of y

OpenStudy (anonymous):

-2

OpenStudy (campbell_st):

thats correct... so on the number plane plot your 3 points (-3, -2) (-1 , 3) (1, 8) if you have calculated correctly you can draw a straight line that passes through all 3 points

OpenStudy (anonymous):

thanks so much that that was correct

OpenStudy (campbell_st):

glad to help